OpenStudy (anonymous):

Calculate the balance after five years for the following account. Account balance: $500.00 Interest: 4.95% Type of Account: Compound Interest

OpenStudy (anonymous):

Okay, I don't need the answer. I just need someone to show me how to get my answer because right now I am really confused on what to do to find my answer. Could anyone help me out?

OpenStudy (anonymous):

\[M = P( 1 + i )^n\] M is the final amount including the principal. P is the principal amount. i is the rate of interest per year. n is the number of years invested. Let's say you have $500 to invest at 4.95% for 5 years. \[M = 500 ( 1+ .0495)^5 = 636.62\]

OpenStudy (anonymous):

LAYMAN: STEP BY STEP year 1: Beg a/c bal-500 int.-(500*.0495)=24.75 ending a/c bal. =500+24.75=524.75 (this is next yrs opening balance) Year 2: beg a/c bal.524.75, int-(524.75*.0495)=25.975125 ending a/c bal= 524.75+25.975125=550.725125 year3: beg a/c bal. 550.725125 int-(550.725125*.0495)=27.26089369 ending a/c bal.=550.725125+27.26089369=577.9860187 year 4: beg a/c bal-577.9860187 int=(577.9860187*.0495)=28.61030793 ending a/c bal=606.5963266 year 5: beg a/c bal=606.5963266 int=(606.5963266*.0495)=30.02651817 ENDING BALANCE=636.6228448